10        20        30        40           50 1pedA T0112  ......MKGFAMLGINKLGWIEKERPVAGSYDAIVRPLAVSPCTSDIHTVFEGA...LGD...RKNMILG masdnl--SAVLYKQNDLRLEQRPIPEPKEDEVLLQMAYVGICGSDVHYYEHGRiad---fivKDPMVIG 60        70        80        90       100       110       120 1pedA T0112  HEAVGEVVEVGSEVKDFKPGDRVIVPCTTPDWRSLEVQAGFQQHSNGMLAGWKFSNFKDGVFGEYFHVND HEASGTVVKVGKNVKHLKKGDRVAVEPGVPCRRCQFCKEGKYNLCPDLTFCA--TPPDDGNLARYYVHAA 130       140       150       160       170       180       190 1pedA T0112  ADMNLAILPKDMPLENAVMITDMMTTGFHGAELADIQMGSSVVVIGIGAVGLMGIAGAKLRGAGRIIGVG D--FCHKLPDNVSLEEGA-LLEPLSVGVHACRRAGVQLGTTVLVIGAGPIGLVSVLAAKAYGAF-VVCTA 200       210         220          230       240       250       260 1pedA T0112  SRPICVEAAKFYGA..TDILNYKNG...HIVDQVMKLTNGKGVDRVIMAGGGSETLSQAVSMVKPGGIIS RSPRRLEVAKNCGAdvTLVVDPAKEeesSIIERIRSAIGD-LPNVTIDCSGNEKCITIGINITRTGGTLM 270       280       290       300       310       320       330 1pedA T0112  NINYHGSGDALLIPRVEWGCGMAHKTIKGGLCPGGRLRAEMLRDMVVYNRVDLSKLVTHVYHGFDHIEEA LVGMGSQ----MVTVPLVNACAREIDIKSVFRY--CNDYPIALEMVASGRCNVKQLVTHS-FKLEQTVDA 340       350 1pedA T0112  LLLMKDKPKDLIKAVVI-l... FEAARKKADNTIKVMIS-crqg

Figure 1: Comparison of candidate ../1pedA-T0112-fssp-global.pw.a2m (lines) to reference ../1pedA-T0112-local.pw.a2m (letters). Reference alignment has 331 matching columns, Candidate alignment has 308 matching columns. 4 were used (for template shift+target shift) to compute shift score. The shift score is 0.000000, with a min shift of 270, max shift of 0, and mean of 0.000000. Optimal subalignment has 270 matching columns, with 313 paired matches. 0 were used (for template shift+target shift) to compute shift score. Optimum shift score is 0.861220, with a min shift of 589, max shift of 0, and mean of 0.193519. Optimum achieved by removing the 81 columns indicated with dotted lines from the candidate alignment.